Your friend in mechanical engineering has invented a money machine. The main drawback of the machine is that it is slow. It takes one year to manufacture . However, once built, the machine will last forever and will require no maintenance. The machine can be built immediately, but it will cost to build. Your friend wants to know if he should invest the money to construct it. If the interest rate is per year, what should your friend do? What is your advice if the machine takes one year to build?
Answer #1

step 1) 800/ 0.30= 26,666666667 or 26,667

step 2) 26,667- 26,050= 617 NPV answer


what should your friend so?

answer: accept the machine because the npv is equal to or greater than $0.


part 2) what is your advise if the machine takes 1 year to build?


step 1)  pv= 1/ 1+0.030 * 800/ 0.030

              0.9708737864 * 26,666666667

             pv=25,889.967637334 or 25,890 


step 2) NPV= 25,890- 26,050= -160 Answer


what should your friend do?

answer: reject the machine because the NPV is less than $0.

answered by: Andrew San Andres